# SemFlow: Binding Semantic Segmentation and Image Synthesis via Rectified Flow

### v1

@@ Abstract
Semantic segmentation and semantic image synthesis are two representative tasks in visual perception and generation. While existing methods consider them as two distinct tasks, we propose a unified diffusion-based framework (SemFlow) and model them as a pair of reverse problems. Specifically, motivated by rectified flow theory, we train an ordinary differential equation (ODE) model to transport between the distributions of real images and semantic masks. As the training object is symmetric, samples belonging to the two distributions, images and semantic masks, can be effortlessly transferred reversibly. For semantic segmentation, our approach solves the contradiction between the randomness of diffusion outputs and the uniqueness of segmentation results. For image synthesis, we propose a finite perturbation approach to enhance the diversity of generated results without changing the semantic categories. Experiments show that our SemFlow achieves competitive results on semantic segmentation and semantic image synthesis tasks. We hope this simple framework will motivate people to rethink the unification of low-level and high-level vision. Project page: https://github.com/wang-chaoyang/SemFlow.
@@ 1 Introduction
Understanding semantic content and creating images from semantic conditions are fundamental research topics in computer vision. Semantic segmentation [48, 7, 62, 87, 11, 10] and image synthesis [8, 31, 26, 74, 91] are two representative dense prediction tasks and inspires various downstream applications, including autonomous driving [5] and medical image analysis [90]. The former aims to assign a category label to each pixel in the image, while the latter aims to generate realistic images given semantic layouts.
Although semantic segmentation and image synthesis constitute a pair of reverse problems, existing works typically solve them using two distinct methodologies. On the one hand, semantic segmentation models mostly follow the spirit of discriminative models. Specifically, a well-pre-trained backbone is employed to extract multi-scale features, and then a task-specific decoder is used for dense prediction. On the other hand, semantic image synthesis frameworks are mainly built upon generative adversarial networks (GAN) [17, 55, 60, 65] or diffusion models (DM) [66, 67, 23]. GAN-based methods [91, 74, 26] typically take semantic layouts as inputs and adopt a discriminator for adversarial training. Meanwhile, DM-based methods [39] generate from noise with a semantic layout functioning as a control signal.
An intuitive solution is to represent the segmentation mask as a colormap and model it as a conditional image generation task [59]. However, there are several implementation challenges. Overall, most of the discriminative segmentation models do not apply to this problem due to their irreversibility. For generative adversarial networks, their generators are typically unidirectional. Jointly training multiple unidirectional models to achieve bidirectional generation [91] is not the concern of this paper. Latent diffusion models (LDMs) have recently demonstrated great potential in generative tasks. Beyond generative tasks, several works [59, 71] attempt to apply diffusion models for segmentation, with images functioning as conditions, but their applicable tasks are limited to be class-agnostic. There are three main problems for existing LDM-based segmentation frameworks: 1) The contradiction between the randomness of generation outputs and the certainty of segmentation labels. 2) The huge cost brought by multiple inference steps. 3) The irreversibility between semantic masks and images.
@@
In this paper, we use rectified flow [43, 45, 15] to enable LDM as a unified framework for semantic segmentation and semantic image synthesis. Our key idea is summarized in Fig. 1. Starting from an LDM [61] framework, we solve the above problems with three methodologies. First, we redefine the mapping function to address the randomness. Previous works [59] aim to learn the mapping from the joint distribution of Gaussian noise and images to the segmentation masks. We argue that Gaussian noise in this mapping function is redundant and negatively affects the determinism of semantic segmentation results. Instead, our model directly learns the mapping from images to masks. Secondly, we make the mapping reversible via rectified flow. Rectified flow is an ordinary differential equation (ODE) framework with a time-symmetric training object. This feature allows the model to be trained once to obtain bi-directional transmission capabilities and can be solved numerically using simple ODE solvers such as Euler. Finally, we propose a finite perturbation method to enable multi-modal image synthesis, as the mapping is one-to-one in semantic segmentation but one-to-many in semantic synthesis. We add amplitude-limited noise, enabling masks to be sampled from a collection of semantic-invariant distributions rather than a fixed value, and further improving the quality of synthesized results. Moreover, our model needs fewer inference steps than traditional diffusion models because the transport trajectory is straight, significantly reducing the gap between LDM and traditional discriminative models in segmentation.
The main contributions are as follows: 1) We introduce SemFlow, a new unified framework that binds semantic segmentation and image synthesis with rectified flow, effectively leveraging the length of the generative models. 2) We propose specialized designs of SemFlow, including pseudo mask modeling, bi-directional training of segmentation and generation, and a finite perturbation strategy. 3) We validate SemFlow on several popular benchmarks. For semantic segmentation, SemFlow greatly narrows the gap between diffusion models and discriminative models in terms of accuracy and inference speed with a more elegant framework. Meanwhile, SemFlow also performs decently in semantic image synthesis tasks.
@@ 2 Related Work
Diffusion Models and Rectified Flow. Diffusion models [23, 66, 67] have shown impressive results in the field of generation, such as image generation [61, 58, 85, 25, 13, 57], video generation [82], image editing [4, 49, 63, 54, 21], image super resolution [64, 24] and point cloud [50, 53, 89, 84, 51]. Most of these methods are based on stochastic differential equations (SDEs) and need multiple steps for generation. Recently, some works [43, 42, 41, 1, 20] propose to model with probability flow ordinary differential equations (ODEs) to reduce the inference steps. Specifically, rectified flow [43, 45] defines the forward process as straight paths and uses reflow to minimize the sampling steps to one. Although rectified flow has demonstrated decent results on image generation and image transfer tasks, they are limited to low-level vision and lack an exploration of the unification of segmentation and generation tasks.
@@
Semantic Segmentation is one of the core tasks in visual perception, aiming to assign each pixel of the given image a semantic category. Previous semantic segmentation approaches are typically built upon discriminative modeling, consisting of a strong backbone [19, 14, 46, 47] for feature extraction and a task-specific decoder head [7, 62, 87, 11, 10, 35, 32, 37, 36, 34] for mask prediction. Recently, some works [33, 88, 16, 79, 77, 3, 72, 9, 18, 2, 28, 78, 73] exploit diffusion models for segmentation. They typically follow the spirit of discriminative models and employ the diffusion model as a feature extractor. Although UniGS [59] and LDMSeg [71] attempt to use a plain Stable Diffusion framework for segmentation, their applicable tasks are limited to be class-agnostic. In practice, reconciling the stochastic outputs of diffusion models with the deterministic results of semantic segmentation is difficult. We rethink this problem and re-model it with rectified flow.
@@
Semantic Image Synthesis is the reverse problem of semantic segmentation, aiming to generate realistic images given semantic layouts [8, 56, 44, 76, 92, 68, 69, 81, 31, 38, 52]. Several studies have delved into semantic synthesis through two methodologies. One methodology [91, 26, 74, 6, 70] is based on GAN and trained with adversarial loss and reconstruction loss. However, some of these methods can only generate unimodal outputs. Although some methods [69, 93] have been developed to address the diversity issues, GAN-based frameworks typically suffer from unstable training and need careful parameter tuning. Another methodology [75] employs diffusion models and regards it as a conditional image generation task, where semantic masks function as control signals. Some works further add a greater variety of control signals to enhance the consistency of synthesized results, like textual prompts [80] and bounding box [39]. Despite these methods achieving good synthetic results, their architecture is usually asymmetric, and the generator is usually unidirectional. This hinders the exploration of the unification of semantic segmentation and image synthesis.
@@ 3 Method
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\alpha_{t}}x_{0},(1-\alpha_{t})I),$ (1)
@@
$\mathcal{L}=\mathbb{E}_{x_{0}\sim q(x_{0}),\epsilon\sim\mathcal{N}(0,I),t}\left[\left|\left|\epsilon_{\theta}(\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t},t)-\epsilon_{t}\right|\right|^{2}\right],$ (2)
@@
Eq. 6 formulates the existing diffusion-based segmentation models with generative modeling. However, this modeling approach suffers from the following problems: 1) The randomness of initial noise and the determinism of the segmentation mask are contradictory. 2) From a transmission point of view, the transfer from noise to the segmentation mask does not conform to the paradigm of semantic segmentation task, where noise is redundant. 3) The images in this approach only serve as the condition, which is non-casual in reverse transfer. Thus, the reversed transfer is a separate problem.
@@ 3.2 Unify Segmentation and Synthesis with Rectified Flow
$m_{0}=s*\lfloor c/k^{2}\rfloor,\quad m_{1}=s*\lfloor(c-m_{0}*k^{2})/k\rfloor,\quad m_{2}=s*(c-m_{0}*k^{2}-m_{1}*k),$ (7)
@@
Note that we do not use image captions or image features as prompts. As our model unifies semantic segmentation and image synthesis with one model, the usage of captions contradicts the definition of semantic segmentation task while the latter is non-casual for image synthesis. Thus, in this work, we set the prompt as empty.
@@
When $\varphi$ is the trajectory of rectified flow [43], $z_{t}$ can be reformulated as the linear interpolation between $z_{0}$ and $z_{1}$ as $z_{t}=(1-t)z_{0}+tz_{1}$. We aim to learn the velocity field using neural networks $v_{\theta}(z_{t},t)$ and solve it with optimization methods,
@@ 4.1 Experimental Settings
Dataset and Metrics. We study SemFlow using three popular datasets: COCO [40], CelebAMask-HQ [29], and Cityscapes [12]. They contain 133, 19, and 19 categories, respectively. The train/val splits follow the setting of these datasets. For semantic segmentation, we evaluate with mean intersection over union (mIoU). For semantic image synthesis, we assess with the Fréchet inception distance (FID) [22] and learned perceptual image patch similarity (LPIPS) [86]. We also evaluate the quality of synthesized images by calculating mIoU with DRN-D-105 [83] model.
Implementation Details. The SemFlow model is built upon Stable Diffusion UNet and initialized with weights from the pre-trained SD 1.5. Images and semantic masks in COCO, CelebAMask-HQ, and Cityscapes are resized and cropped into $512\times 512$, $512\times 512$, and $512\times 1024$, and trained for 80K, 80K, and 8K iterations, respectively. The batch size is 256. We use the off-the-shelf VAE of the corresponding Stable Diffusion model. The spacing $s$ is set as 50, and the division coefficient $k$ is 6. We use the AdamW optimizer with $1\times 10^{-3}$ weight decay. The initial learning rate is $1\times 10^{-4}$, $5\times 10^{-5}$, and $2\times 10^{-5}$ for COCO, Cityscapes, and CelebAMask-HQ, respectively. All experiments are conducted on 8 NVIDIA A100 GPUs. Note that there is no need to train two models for segmentation and synthesis separately since the optimization target in Eq. 10 is time-symmetric.
Baselines. We seek to unify semantic segmentation (SS) and semantic image synthesis (SIS) into a pair of reverse problems and train one model for a solution. As few previous works achieve this goal, we thus take different baselines for SS and SIS, respectively. For SS, we design two baseline models, DSMs [61], which follow the principle of diffusion-based conditional generation modeling in Sec. 3.1. The network structure and hyperparameters of DSM follow SemFlow, except that the inputs of UNet are 8 channels. Specifically, the noise and images are concatenated in the channel dimension. DSM-DDIM and DSM-DDPM differ in differential equation modeling (ODE vs. SDE). For SIS, we take pix2pixHD [74], SPADE [56], SC-GAN [76], BBDM [30] and CycleGAN [91] as the baselines.
@@ 4.2 Main Results
[…]
Method Task Sampler mIoU
SS SIS
Mask2Former [10] - 61.5
DSM [61] DDIM-50 23.1
DSM [61] DDPM-50 27.5
SemFlow (ours) Euler-10 49.2
Figure 2: Semantic segmentation results on COCO dataset. For the ground truth, each color reflects the value of anchors (Eq. 7), which corresponds to one semantic category, and the color white indicates the ignored regions. The predictions of DSM vary considerably under different random seeds.
@@
Tab. 1 shows the quantitative results on the COCO dataset. We compare SemFlow with two variants of DSM and choose Mask2Former as the upper bound, which is a strong discriminative segmentation model. Regarding accuracy, our SemFlow achieves 49.2 mIoU and outperforms DSMs by a remarkable margin. Moreover, SemFlow only uses a simple sampler, forward Euler method, and fewer inference steps than DSMs. Further reducing the inference steps brings about faster generation but witnesses a slight drop in performance, which is analyzed in Sec. 4.3.
@@
Table 2: Quantitative results for semantic image synthesis. SS and SIS represents semantic segmentation and semantic image synthesis, respectively.
Method Task CelebAMask-HQ Cityscapes
SS SIS FID LPIPS mIoU FID
pix2pixHD [74] 54.7 0.529 58.3 95.0
SPADE [56] 42.2 0.487 62.3 71.8
SC-GAN [76] 19.2 0.395 66.9 49.5
BBDM [30] 21.4 0.370 N/A N/A
SemFlow (ours) 32.6 0.393 61.8 40.2
Semantic Image Synthesis. We compare our approach with other specialist models on semantic image synthesis tasks in Tab. 2. On CelebAMask-HQ, SemFlow achieves a performance of 32.6 FID and 0.393 LPIPS. On Cityscapes, our approach improves urban streets synthesis to FID of 40.2, compared to FID of SC-GAN 49.5.
@@
Discussion. Although the performance gap exists compared with Mask2Former, our approach solves the two major issues of diffusion-based segmentation models: 1) Randomness of outputs and 2) large inference steps. We believe it significantly reduces the gap between discriminative models and diffusion models.

### v2

@@ Abstract
footnotetext: Corresponding author: Lu Qi and Xiangtai Li.
Semantic segmentation and semantic image synthesis are two representative tasks in visual perception and generation. While existing methods consider them as two distinct tasks, we propose a unified framework (SemFlow) and model them as a pair of reverse problems. Specifically, motivated by rectified flow theory, we train an ordinary differential equation (ODE) model to transport between the distributions of real images and semantic masks. As the training object is symmetric, samples belonging to the two distributions, images and semantic masks, can be effortlessly transferred reversibly. For semantic segmentation, our approach solves the contradiction between the randomness of diffusion outputs and the uniqueness of segmentation results. For image synthesis, we propose a finite perturbation approach to enhance the diversity of generated results without changing the semantic categories. Experiments show that our SemFlow achieves competitive results on semantic segmentation and semantic image synthesis tasks. We hope this simple framework will motivate people to rethink the unification of low-level and high-level vision.
@@ 1 Introduction
Understanding semantic content and creating images from semantic conditions are fundamental research topics in computer vision. Semantic segmentation [45; 7; 59; 85; 11; 10] and image synthesis [8; 32; 28; 72; 90] are two representative dense prediction tasks and inspires various downstream applications, including autonomous driving [5] and medical image analysis [89]. The former aims to assign a category label to each pixel in the image, while the latter aims to generate realistic images given semantic layouts.
Although semantic segmentation and image synthesis constitute a pair of reverse problems, existing works typically solve them using two distinct methodologies. On the one hand, segmentation models mostly follow the spirit of discriminative models. Specifically, a pre-trained backbone is employed to extract multi-scale features, and then task-specific decoders are used for dense prediction. On the other hand, semantic image synthesis frameworks are mainly built upon generative adversarial networks (GAN) [19; 52; 57; 62] or diffusion models (DM) [63; 64; 25]. GAN-based methods [90; 72; 28] typically take semantic layouts as inputs and adopt a discriminator for adversarial training. Meanwhile, DM-based methods [36] generate from noise with semantic layouts functioning as control signals.
An intuitive solution is to represent the segmentation mask as a colormap and model it as a conditional image generation task [56]. However, there are several implementation challenges. Overall, most of the discriminative segmentation models do not apply to this problem due to their irreversibility. For generative adversarial networks, their generators are typically unidirectional. Jointly training multiple unidirectional models to achieve bidirectional generation [90] is not the concern of this paper. Latent diffusion models (LDMs) have recently demonstrated great potential in generative tasks. Beyond generative tasks, several works [56; 69] attempt to apply diffusion models for segmentation, with images functioning as conditions, but their applicable tasks are limited to be class-agnostic. There are three main problems for existing LDM-based segmentation frameworks: 1) The contradiction between the randomness of generation outputs and the certainty of segmentation labels. 2) The huge cost brought by multiple inference steps. 3) The irreversibility between semantic masks and images.
@@
In this paper, we use rectified flow [40; 42; 17] to enable LDM as a unified framework for semantic segmentation and semantic image synthesis. Our key idea is summarized in Fig. 1. Starting from an LDM [58] framework, we solve the above problems with three methodologies. First, we redefine the mapping function to address the randomness. Previous works [56] aim to learn the mapping from the joint distribution of Gaussian noise and images to the segmentation masks. We argue that Gaussian noise in this mapping function is redundant and negatively affects the determinism of semantic segmentation results. Instead, our model directly learns the mapping from images to masks. Secondly, we make the mapping reversible via rectified flow. Rectified flow is an ordinary differential equation (ODE) framework with a time-symmetric training object. This feature allows the model to be trained once to obtain bi-directional transmission capabilities and can be solved numerically using simple ODE solvers such as Euler. Finally, we propose a finite perturbation method to enable multi-modal image synthesis, as the mapping is one-to-one in semantic segmentation but one-to-many in semantic synthesis. We add amplitude-limited noise, enabling masks to be sampled from a collection of semantic-invariant distributions rather than a fixed value, and further improving the quality of synthesized results. Moreover, our model needs fewer inference steps than traditional diffusion models because the transport trajectory is straight, significantly reducing the gap between LDM and traditional discriminative models in segmentation.
The main contributions are as follows: 1) We introduce SemFlow, a new unified framework that binds semantic segmentation and image synthesis with rectified flow, effectively leveraging the length of the generative models. 2) We propose specialized designs of SemFlow, including pseudo mask modeling, bi-directional training of segmentation and generation, and a finite perturbation strategy. 3) We validate SemFlow on several popular benchmarks. For semantic segmentation, SemFlow dramatically narrows the gap between diffusion models and discriminative models in terms of accuracy and inference speed with a more elegant framework. Meanwhile, SemFlow also performs decently in semantic image synthesis tasks.
@@ 2 Related Work
Diffusion Models and Rectified Flow. Diffusion models [25; 63; 64] have shown impressive results in the field of generation, such as image generation [58; 55; 83; 27; 13; 54], video generation [80], image editing [4; 46; 60; 51; 23], image super resolution [61; 26] and point cloud [47; 50; 88; 82; 48]. Most of these methods are based on stochastic differential equations (SDEs) and need multiple steps for generation. Recently, some works [40; 39; 38; 1; 22] propose to model with probability flow ordinary differential equations (ODEs) to reduce the inference steps. Specifically, rectified flow [40; 42] defines the forward process as straight paths and uses reflow to minimize the sampling steps to one. Although rectified flow has demonstrated decent results on image generation and image transfer tasks, they are limited to low-level vision and lack an exploration of the unification of segmentation and generation tasks.
@@
Semantic Segmentation is one of the core tasks in visual perception, aiming to assign each pixel of the given image a semantic category. Previous semantic segmentation approaches are typically built upon discriminative modeling, consisting of a strong backbone [21; 16; 43; 44; 33] for feature extraction and a task-specific decoder head [7; 59; 85; 11; 10; 87; 65; 81; 15; 14] for mask prediction. Recently, some works [34; 86; 18; 77; 75; 3; 70; 9; 20; 2; 29; 76; 71] exploit diffusion models for segmentation. They typically follow the spirit of discriminative models and employ the diffusion model as a feature extractor. Although UniGS [56] and LDMSeg [69] attempt to use a plain Stable Diffusion framework for segmentation, their applicable tasks are limited to be class-agnostic. In practice, reconciling the stochastic outputs of diffusion models with the deterministic results of semantic segmentation is difficult. We rethink this problem and re-model it with rectified flow.
@@
Semantic Image Synthesis is the reverse problem of semantic segmentation, aiming to generate realistic images given semantic layouts [8; 53; 41; 74; 91; 66; 67; 79; 32; 35; 49]. Several studies have delved into semantic synthesis through two methodologies. One methodology [90; 28; 72; 6; 68] is based on GAN and trained with adversarial loss and reconstruction loss. However, some of these methods can only generate unimodal outputs. Although some methods [67; 92] have been developed to address the diversity issues, GAN-based frameworks typically suffer from unstable training and need careful parameter tuning. Another methodology [73] employs diffusion models and regards it as a conditional image generation task, where semantic masks function as control signals. Some works further add a greater variety of control signals to enhance the consistency of synthesized results, like textual prompts [78] and bounding box [36]. Despite these methods achieves good synthetic results, their architecture is usually asymmetric, and the generator is usually unidirectional. This hinders the exploration of the unification of semantic segmentation and image synthesis.
@@ 3 Method
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\alpha_{t}}x_{0},(1-\alpha_{t})I),\quad t\in[0..T],$ (1)
@@
$\mathcal{L}=\mathbb{E}_{x_{0}\sim q(x_{0}),\epsilon\sim\mathcal{N}(0,I),t}\left[\left|\left|\epsilon_{\theta}(\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon_{t})-\epsilon_{t}\right|\right|_{2}^{2}\right],$ (2)
@@
Eq. 6 formulates the existing diffusion-based segmentation models with generative modeling. However, this modeling approach suffers from the following problems: 1) The randomness of initial noise and the determinism of the segmentation mask are contradictory. 2) From a transmission point of view, the transfer from noise to the segmentation mask does not conform to the paradigm of semantic segmentation task, where noise is redundant. 3) The images in this approach only serve as the condition, which is non-causal in reverse transfer. Thus, the reversed transfer is a separate problem.
@@ 3.2 Unify Segmentation and Synthesis with Rectified Flow
$m_{0}^{\prime}=\lfloor c/k^{2}\rfloor,\;m_{1}^{\prime}=\lfloor(c-m_{0}^{\prime}*k^{2})/k\rfloor,\;m_{2}^{\prime}=c-m_{0}^{\prime}*k^{2}-m_{1}^{\prime}*k,\;M=s*(m_{0}^{\prime},m_{1}^{\prime},m_{2}^{\prime}),$ (7)
@@
Note that we do not use image captions or image features as prompts. As our model unifies semantic segmentation and image synthesis with one model, the usage of captions contradicts the definition of semantic segmentation task while the features are non-causal for image synthesis. Thus, in this work, we set the prompt as empty.
@@
When $\varphi$ is the trajectory of rectified flow [40], $z_{t}$ can be reformulated as the linear interpolation process between $z_{0}$ and $z_{1}$ as $z_{t}=(1-t)z_{0}+tz_{1}$. We aim to learn the velocity field using neural networks $v_{\theta}(z_{t},t)$ and solve it with optimization methods,
@@ 4.1 Experimental Settings
Dataset and Metrics. We study SemFlow using three popular datasets: COCO-Stuff [37], CelebAMask-HQ [30], and Cityscapes [12]. They contain 171, 19, and 19 categories, respectively. For semantic segmentation, we evaluate with mean intersection over union (mIoU). For semantic image synthesis, we assess with the Fréchet inception distance (FID) [24] and learned perceptual image patch similarity (LPIPS) [84].
Implementation Details. The SemFlow model is built upon Stable Diffusion UNet and initialized with weights from the pre-trained SD 1.5. Images and semantic masks in COCO-Stuff, CelebAMask-HQ, and Cityscapes are resized and cropped into $512\times 512$, $512\times 512$, and $512\times 1024$, respectively. We use the off-the-shelf VAE of the corresponding Stable Diffusion model. The spacing $s$ is set as 50, and the division coefficient $k$ is 6. The amplitude of perturbation is 6. Note that there is no need to train two models for segmentation and synthesis separately since the optimization target in Eq. 10 is time-symmetric.
Baselines. We seek to unify semantic segmentation (SS) and semantic image synthesis (SIS) into a pair of reverse problems and train one model for a solution. As few previous works achieve this goal, we thus take different baselines for SS and SIS, respectively. For SS, we design two baseline models, DSMs [58], which follow the principle of diffusion-based conditional generation modeling in Sec. 3.1. The network structure and hyperparameters of DSM follow SemFlow, except that the inputs of UNet are eight channels. Specifically, the noise and images are concatenated in the channel dimension. DSM-DDIM and DSM-DDPM differ in differential equation modeling (ODE vs. SDE). For SIS, we take pix2pixHD [72], SPADE [53], SC-GAN [74], BBDM [31] and CycleGAN [90] as the baselines.
@@ 4.2 Main Results
Table 1: Semantic segmentation results on COCO-Stuff dataset. SS and SIS represents semantic segmentation and semantic image synthesis, respectively. Sampler-N means the usage of a specific sampler with N inference steps.
Method Task Sampler COCO-Stuff CelebAMask-HQ
SS SIS mIoU (SS) FID (SIS) LPIPS (SIS)
MaskFormer [11] - 41.9 - -
DSM [58] DDIM-200 16.1 - -
DSM [58] DDPM-200 20.2 - -
pix2pixHD [72] - - 54.7 0.529
SPADE [53] - - 42.2 0.487
SC-GAN [74] - - 19.2 0.395
BBDM [31] BBDM-200 - 21.4 0.370
SemFlow (ours) Euler-25 38.6 32.6 0.393
Figure 2: Semantic segmentation results on COCO-Stuff dataset. For the ground truth, each color reflects the value of anchors (Eq. 7), which corresponds to one semantic category, and the color white indicates the ignored regions. The predictions of DSM vary considerably under different random seeds.
@@
Tab. 1 shows the quantitative results on the COCO-Stuff dataset. We compare SemFlow with two variants of DSM and MaskFormer, which is a classical discriminative segmentation model. Regarding accuracy, our SemFlow achieves 38.6 mIoU and outperforms DSMs by a remarkable margin. Moreover, SemFlow only uses a simple sampler, forward Euler method, and fewer inference steps than DSMs. Further reducing the inference steps brings about faster generation but witnesses a slight drop in performance, which is analyzed in Sec. 4.3.
@@
Semantic Image Synthesis. We compare our approach with other specialist models on semantic image synthesis tasks in Tab. 1. On CelebAMask-HQ, SemFlow achieves a performance of 32.6 FID and 0.393 LPIPS.
@@
Discussion. Although the performance gap exists compared with specialist models, our approach solves the two major issues of diffusion-based segmentation approaches: 1) Randomness of outputs and 2) Large inference steps. We believe it significantly reduces the gap between discriminative models and diffusion models.